Ask your own question, for FREE!
Mathematics 13 Online
OpenStudy (anonymous):

From the deck of cards shown, two cards are drawn with replacement. The probability of drawing a yellow followed by a yellow is 1 out of ____.

OpenStudy (anonymous):

|dw:1396242169344:dw|

OpenStudy (raden):

P(Y and Y) = P(Y) * P(Y|Y) = 1/10 * 1/10 = 1/... ?

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!